OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the x's to add out? 2x + 3y = 8 -6x + 4y = 11 3 -3 6 none of the above Can someone show me how to solve this? ..please?

OpenStudy (ranga):

There is a -6x in the second equation. To eliminate -6x we need to add +6x to it. If you multiply the first equation by 3 and then add it to the second equation then the x term will be eliminated.

OpenStudy (anonymous):

Thank you that's what ive been trying to figure out for about 30 minutes

OpenStudy (ranga):

2x + 3y = 8 multiply by 3 6x + 9y = 24 add it to the second equation: 6x + 9y = 24 -6x + 4y = 11 add 0x + 13y = 35 13y = 35 y = 35/13 They are not asking you find x and y. But I am just showing how to do it for y. You can put the y value in either of the equations and then solve for x.

OpenStudy (anonymous):

That helped a lot, thank you

OpenStudy (ranga):

you are welcome.